OpenStudy (lilmike234):

what is the radius of a circle with the equation x^2+y-3^2 =21

OpenStudy (ac3):

The formula for a circle is x^2+y^2=r^2, therefore it's already set up for you. Not it's not 21 Here I'm gonna go with the sqrt(21) for the radius. Unless 21 is supposed to be squared in that case it's just 21.

OpenStudy (zzr0ck3r):

\(x^2+(y-3)^2=21\\ (x+0)^2+(y-3)^2=(\sqrt{21})^2\)

OpenStudy (zzr0ck3r):

You should be able to complete it from here?

OpenStudy (wolf1728):

Is that equation written exactly as the problem states or should it be: x^2+(y-3)^2 =21

OpenStudy (lilmike234):

So my answer is 4.583

OpenStudy (zzr0ck3r):

How did you get that?